Systems Staff Meeting Notes 8/21/00

Systems Staff Meeting Notes
August 21, 2000

Present: Rich Bennett, Cynthia Bowker, Will Chaney, Bill Covey, Debra Harris, Winston Harris, Martha Hruska, Bobbie Meng, Suzy Shaw, Hilton Soto, Geoffrey Stewart, Ying Tang
Announcements
  1. Martha will be working in Systems in the morning hours beginning Wednesday, August 23 at 8:30 AM.
  2. The auto-feed on the copier is failing. Debra will call for service.
  3. Rich Bennett's father celebrated his 90th birthday yesterday.
  4. Hilton will be getting engaged soon - he hasn't asked her yet, but will as soon as he gets the ring.
  5. Two students came to the Systems front door looking for the "Library Computer Lab". They weren't looking for the laptop plug-in area; they were looking for free computers to use. Bill sent them over to West and notes that this is how the library computers are being advertised.
  6. Will will handle the Liaison Meetings (first Thursday of every month).
Review Minutes of 8/7/00
  1. Re: peripherals: #2 Announcements - Bill notes that there are now monitors that are manufactured to fold up and fit into your pocket. It's like an electronic book.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. Computer replacement orders will resume in the near future. No word on when FCLA funds will be released, but it should be soon. We can begin to assemble orders - 20 workstations at a time. Martha will inquire further in the next two weeks as to which machines need to be upgraded. We have a lot of machines that will need HD or memory upgrades - would be better to just replace them. The Baldwin grant will provide two new positions and they will need new machines.
  2. Distribution list for Visors - we can now order green & orange. Martha needs to get a list of the people who have already received one so she can make sure all of the department chairs have had the option to get one before she makes a new distribution list.
  3. A machine for Cynthia has been ordered. It should be shipped within 5-7 days. Suzy will continue to work on vacating room 18A.
  4. New Room 2 (surveyed equipment) procedures - Systems is no longer maintaining the inventory. We'll meet the liaison returning the old equipment in room 2 and note what they've turned in. After deleting it from the IR database, we'll notify Bill Hanssen, who will be responsible for surveying it from that point.
Networking
  1. The fall crunch is on this week and next. Expect the Internet POP to get slow down. The semester startup always means heavier traffic. The NERDC Network Service page (http://net-services.ufl.edu/) shows a graph of inbound/outbound UF network traffic. When it hits 100%, that's when there's trouble. Within campus, traffic should be fine. It's accessing the off-campus sites that become problematic. The mainframe, Dean's Menu and CICS will choke; NOTIS will be relatively unaffected.
  2. Cataloger's Desktop/Plus - the license had to be renewed, had some problems fixing that. Once that was solved, there was one user who was unable to get in. Will finally got that problem fixed. This software requires a quarterly update - but the license is good for six months.
  3. 2nd & 3rd wiring - project is hold at the moment. Rob Roberts ran out of molding, which stopped Digital. Things should resume tomorrow. The wires pulled on the second floor have not been punched down. They are still sharing phones on these floors.
  4. Wireless - money is available, waiting until Brad has time to do the electrical - then Digital will need to run the wires in the stack. Another access point needs to be ordered.
  5. VISIO Wiring Documentation - Suzy has been working on this - it goes as far as the switches and servers. The web server for the switches is not real robust; it could overload. We may make this available to super liaisons, but not to the general public. It will be used as a quick reference tool - hopefully it will be easily updated.
  6. Network wiring adjustments - awaiting an estimate from Judy on the West classroom, room 124 closet and the branches. Will wait for completion on the 2nd & 3rd floors before working on the North tower.
Software Projects Status
  1. Film Log - new batch of changes on the Q&A replacement database. Hilton will adjust; then they'll be ready for the data transfer.
  2. The Employee-Patron update ran this week - all problems have been fixed.
  3. The Student-Patron update ran okay.
  4. IR Database - Winston is working on 3 features: creating a read-only, web accessible version for Facilities, label writing integration and some minor updates for Bobbie.
DP Coordinator
  1. Students will return on Wednesday.
  2. Debra turned in the annual Library Statistics last week.
Calendar
  1. Debra will be out tomorrow morning.
  2. Labor Day holiday - September 4.
  3. Martha will be out September 11, 26 & 27.
  4. LMS Conference in Orlando - Oct. 2-4 (Martha will be out).


Back to the Systems Homepage

Last updated August 28, 2000
by Debra Harris
debharr@mail.uflib.ufl.edu